You are buying flowers to hand out at a school dance. Roses Cost $30 for a dozen but cost more if bought individually. With the money you have you can buy 7 dozens and 4 single roses, or 64 single roses. How Much Is one rose? How Much Money do you have?
Answer by amazingrace333(11)

What do you know...you know that:
Cost of one dozen roses = $30
You can buy 7 dozen roses + 4 single roses = cost of 64 single roses
You want to know the cost of a single rose and the total amount of money you have.
If X = the cost of a single rose and $30 = the cost of a dozen roses, then your equation is:
7($30) + 4X = 64X
Simply and solve for X.
$210 + 4X = 64X
$210 + (4X - 4X) = 64X - 4X
$210 = 60X
$210/60 = 60X/60
$3.50 = X
Therefore, the cost of a single rose is $3.50.
Now input that cost into the original equation to get the total amount of money you have to spend.
7($30) + 4($3.50) =
$210 + $14
$224
Check your answer:
64 individual roses = 64($3.50) or $224.00
The cost of a single rose is $3.50 and you have $224.00 to spend on roses.